A car is stopped at a red light. the light turns green and the driver accelerates at a speed of 3.5 meters per second squared, for 8 seconds. How far has the car traveled during these 8 seconds?
Answer #1

a = 3.5m/s^2

t = 8s

S = ut + 1/2 at^2

=) S = 0 + 0.5×3.5×8^2 = 112 m.

That means required answer is S = 112m